# Hierarchical Vision-Language Planning for Multi-Step Humanoid Manipulation

### v1

@@ Abstract
Enabling humanoid robots to reliably execute complex multi-step manipulation tasks is crucial for their effective deployment in industrial and household environments. This paper presents a hierarchical planning and control framework designed to achieve reliable multi-step humanoid manipulation. The proposed system comprises three layers: (1) a low-level RL-based controller responsible for tracking whole-body motion targets; (2) a mid-level set of skill policies trained via imitation learning that produce motion targets for different steps of a task; and (3) a high-level vision-language planning module that determines which skills should be executed and also monitors their completion in real-time using pretrained vision-language models (VLMs). Experimental validation is performed on a Unitree G1 humanoid robot executing a non-prehensile pick-and-place task. Over 40 real-world trials, the hierarchical system achieved a 72.5% success rate in completing the full manipulation sequence. These experiments confirm the feasibility of the proposed hierarchical system, highlighting the benefits of VLM-based skill planning and monitoring for multi-step manipulation scenarios. See https://vlp-humanoid.github.io/ for video demonstrations of the policy rollout.
@@ I Introduction
We empirically validate our proposed hierarchical system on the 29-DoF Unitree G1 humanoid robot, performing a representative multi-step pick-and-place task. In real-world testing, the integrated system successfully completed the full manipulation sequence in 72.5% of 40 trials. Our hierarchical approach demonstrates effective coordination of manipulation skills and visual feedback, highlighting the benefit of structured skill decomposition and VLM-based execution monitoring for reliable, multi-step humanoid autonomy.
@@ II-A Hierarchical Control in Humanoids
While low-level policies reliably track motion targets, generating these references is equally critical for accomplishing manipulation tasks. H2O (Human-to-Humanoid) [17] utilized single-camera human pose estimation to enable zero-shot teleoperation. OmniH2O [14] extended H2O’s concept, offering a unified interface for teleoperation using various modalities, including RGB cameras, VR, motion capture suits, and exoskeleton systems. OmniH2O further incorporated diffusion-based imitation learning [7], enabling the autonomous acquisition of manipulation skills directly from human demonstrations. Similarly, HumanPlus [11] adopted this hierarchical structure, leveraging human-teleoperated demonstrations recorded along with proprioceptive and egocentric visual data. These demonstrations were used to train the Humanoid Imitation Transformer (HIT)—a transformer-based imitation learning policy inspired by the Action Chunking Transformer (ACT) [31]. Using only 20–40 demonstrations per skill, HumanPlus reported 60–100
@@ III Method
Figure 2: Overview of the proposed hierarchical framework for autonomous multi-step humanoid manipulation. The system consists of: (1) a high-level planning module using vision-language models (VLMs) to generate task-specific skill sequences and monitor their execution via visual feedback; (2) a mid-level motion-generation module composed of imitation-learned action expert policies, mapping egocentric visual inputs and proprioceptive robot states into target joint configurations; and (3) a low-level reinforcement-learning-based tracking policy generating feasible joint-angle trajectories, executed by a high-frequency PD controller.
We propose a three-layer hierarchical control and planning framework for multi-step humanoid manipulation, illustrated in Fig. 2. The system consists of: (1) a low-level whole-body tracking controller trained via reinforcement learning (Sec. III-A), (2) a mid-level motion-generation module composed of multiple imitation-learned action expert policies, each trained from teleoperated demonstrations (Sec. III-B);, and (3) a high-level vision-language-based planning and execution monitoring module (Sec. III-C). The following sections detail each component.
@@ III-B2 Imitation Learning Module
For the IL policies, we adopt the Humanoid Imitation Transformer (HIT) [11], designed specifically for high-DoF humanoids with binocular vision. It is based on the Action Chunking Transformer (ACT) [31] model, which have demonstrated improved robotic control by predicting multi-step future actions in one forward pass. At inference, HIT outputs action chunks of 50 joint-angle targets at 25 Hz. During training, an auxiliary $L_{2}$ loss between predicted and actual future visual embeddings improves visual grounding and generalization. The predicted joint angles are subsequently transformed into corresponding 3D keypoints via forward kinematics, matching the inputs required by the low-level tracking controller.
@@
containing synchronized state-action pairs recorded at 25 Hz. We collect data for the representative task rearrangement involving picking a bag from a table and and placing it onto another. Only successful sequences (30 in total) are retained, segmented into separate skill-specific datasets, i.e. picking and placing, for training two policies. IL policy training employs the demonstration data with hyperparameters specified in Table V. After training, we evaluate independently every skill policy, evaluation over 30 autonomous trials showed success rates of $90\%$ for picking and $83.3\%$ for placing.
@@ III-C3 Skill Library and PDDL-like Representation
We represent each skill available to the robot using structured, human-readable descriptions that include the ‘preconditions’ and ‘effects’ fields of an operator definition in Planning Domain Definition Language (PDDL) [1]. Each skill description includes a one-sentence summary, explicit natural-language preconditions, effects, and example verification queries, as illustrated for the Pick skill in Figure 3. Unlike rigid symbolic representations, this PDDL-like form balances logical rigor and flexibility, enabling the planner to perform common-sense reasoning without exhaustive state enumeration.
@@ III-D Full Hierarchical System
The complete integrated system implements the hierarchical pipeline presented in Figure 2. The pipeline comprises: (1) a high-level VLM Planner (GPT-4o) that generates a sequence of skills from initial visual inputs and natural-language task instructions; (2) a real-time VLM Skill Monitor (Gemini-2.0-Flash-Lite) that verifies skill execution status based on visual feedback; (3) a Skill Selector that dynamically chooses the current skill to execute based on the planned sequence and monitor feedback; (4) a mid-level Imitation Learning Module that activates the corresponding imitation-learned policy according to the selected skill, producing joint-motion targets from egocentric images and proprioceptive data; and (5) a low-level Tracking Policy combined with a high-frequency PD Controller, converting these joint targets into executable motor torque commands.
@@ IV-C Experimental Procedure
A total of 40 independent trials were conducted. The robot operated fully autonomously, executing plans generated by the high-level VLM planner, with skills dynamically selected and monitored for completion in real-time by the VLM execution monitor. Success was defined by the robot correctly picking up the bag, and successfully placing it on the designated target surface.
@@ IV-D Experimental Results
Over 40 trials, the integrated hierarchical system successfully completed the entire manipulation sequence with an overall success rate of $72.5\%$. During each trial, we recorded failures at three distinct levels of the hierarchical framework. The observed failures were categorized as follows, in order of frequency:
@@ V Discussion
The experiments demonstrate that augmenting the canonical two-layer humanoid stack with a vision-language planning and monitoring layer enables reliable execution of multi-step non-prehensile manipulation routines in the physical world. Deployed on a 29-DoF Unitree G1, our proposed three-level architecture successfully completed the obstructed bag-transfer task in 72.5% of forty independent trials. The execution monitor’s one-second mean verification latency at 1Hz fits comfortably within the robot’s motion cycle, and coexists with high-frequency skill policies and RL controller, illustrating the large-model reasoning and perception need not preclude high-rate control.
Compared to prior humanoid manipulation systems, which often rely on open-loop sequences of predefined skills, our approach introduces continuous closed-loop reasoning. By integrating interpretable VLM-based planning with real-time visual verification, our framework dynamically selects appropriate skills, and explicitly determines when transitions between skills should occur.
@@ V-A Failures and Limitations
A qualitative failure analysis reveals three dominant modes. First, skill-policy drift occurs when partial occlusions or out-of-distribution object poses cause the imitation policy to stall short of a grasp or placement. Second, planner mis-grounding emerges when the deliberative VLM hallucinates objects, or fails to sequence skills correctly. This typically involves adding additional irrelevant plan steps, or formulating redundant transition questions. Third, optimistic monitor judgments occasionally trigger a premature transition when subtle geometric conditions — such as a bag resting on the edge of a table — are mis-classified as success.
Several limitations temper the present findings. Our chosen evaluation task family, while non-trivial, is confined to table-top pick-place scenes and therefore under-samples the combinatorial diversity of domestic manipulation. The 1Hz verification rate, although adequate for deliberative motions, would be insufficient for fast skills, such as rapid recovery motions. This implicitly defines the timescale of skill abstraction.
@@ V-B Future work
Error detection and adaptive re-planning:  A critical future enhancement includes adapting the VLM execution monitor system to explicitly detect execution failures or anomalous states. Through failure detection, we can trigger real-time adaptive re-planning. This promises to improve the robustness of the system.
Enhanced semantic and contextual conditioning:  Future work involves incorporating richer conditioning beyond visual and proprioceptive states from the high-level execution monitor. For example, replacing the ResNet [13] backbone with that of CLIP [25] would allow the VLM to have finer-grained semantic control of the skills via text-conditioning.
Expanding the skill library:  Finally, the current skill library is focused primarily on pick-and-place. An immediate next step involves expanding the robot’s skill repertoire. Examples include non-prehensile bi-manipulation and tool use. This broader skill-set would enable more complex, realistic long-horizon tasks, requiring the planning and monitoring components to handle increasingly diverse and dynamic environments. We’ve attempted to include the push skills that are triggered when an obstacle is in the way of the placing skill, as shown in Fig. 1, but we didn’t have time to finish a new set of experiments before submission.
@@ VI Conclusion
In this work, we presented a hierarchical planning and control framework for multi-step humanoid manipulation tasks, combining robust RL-based whole-body tracking, imitation-learned skill policies, and a high-level vision-language planning and monitoring module. Evaluated on the Unitree G1 humanoid robot, our integrated system demonstrated successful autonomous execution of complex manipulation sequences, achieving a 72.5% success rate across real-world trials. Crucially, our approach offers inherent modularity, interpretability, and adaptability—enabling straightforward extension of the robot’s skill repertoire. This work highlights the effectiveness of integrating vision-language models with hierarchical control, paving the way toward increasingly capable autonomous humanoid systems operating in dynamic human environments.

### v3

@@ Abstract
Enabling humanoid robots to reliably execute complex multi-step manipulation tasks is crucial for their effective deployment in industrial and household environments. This paper presents a hierarchical planning and control framework designed to achieve reliable multi-step humanoid manipulation. The proposed system comprises three layers: (1) a low-level controller trained via reinforcement learning responsible for tracking whole-body motion targets, (2) a mid-level set of skill policies trained via imitation learning that produce motion targets for different steps of a task, and (3) a high-level vision-language planning module that determines which skills should be executed and also monitors their completion in real time using pretrained vision-language models (VLMs). Experimental validation is performed on a Unitree G1 humanoid robot executing a non-prehensile pick-and-place task. Over 40 real-world trials, the hierarchical system achieved a 73% success rate in completing the full manipulation sequence. These experiments confirm the feasibility of the proposed hierarchical system, highlighting the benefits of VLM-based skill planning and monitoring for multi-step manipulation scenarios. See https://vlp-humanoid.github.io/ for a video demonstration of the policy rollout.
@@ I Introduction
We empirically validate our proposed hierarchical system on the 29-DoF Unitree G1 humanoid robot, performing a representative multi-step pick-and-place task. In real-world testing, the integrated system successfully completed the full manipulation sequence in 73% of 40 trials. Our hierarchical approach demonstrates effective coordination of manipulation skills and visual feedback, highlighting the benefit of structured skill decomposition and VLM-based execution monitoring for reliable, multi-step humanoid autonomy.
@@ II-A Hierarchical Control in Humanoids
While low-level policies reliably track motion targets, generating these references is equally critical for accomplishing manipulation tasks. H2O (Human-to-Humanoid) [15] utilized single-camera human pose estimation to enable zero-shot teleoperation. OmniH2O [12] extended H2O’s concept, offering a unified interface for teleoperation using various modalities, including RGB cameras, VR, motion capture suits, and exoskeleton systems. OmniH2O further incorporated diffusion-based imitation learning [7], enabling the autonomous acquisition of manipulation skills directly from human demonstrations. Similarly, HumanPlus [10] adopted this hierarchical structure, leveraging human-teleoperated demonstrations recorded along with proprioceptive and egocentric visual data. These demonstrations were used to train the Humanoid Imitation Transformer (HIT)—a transformer-based imitation learning policy inspired by the Action Chunking Transformer (ACT) [31]. Using only 20–40 demonstrations per skill, HumanPlus reported 60–100% success rates across diverse whole-body manipulation tasks.
@@ III Method
Figure 2: Overview of the proposed hierarchical framework for autonomous multi-step humanoid manipulation. The system consists of: (1) a high-level planning module using vision-language models (VLMs) to generate task-specific skill sequences and monitor their execution via visual feedback, (2) a mid-level motion-generation module composed of imitation-learned action expert policies, mapping egocentric visual inputs and proprioceptive robot states into target joint configurations, and (3) a low-level reinforcement-learning-based tracking policy generating feasible joint-angle trajectories, executed by a high-frequency PD controller.
We propose a three-layer hierarchical control and planning framework for multi-step humanoid manipulation, illustrated in Fig. 2. The system consists of: (1) a low-level whole-body tracking controller trained via reinforcement learning (Sec. III-A), (2) a mid-level motion-generation module composed of multiple imitation-learned action expert policies, each trained from teleoperated demonstrations (Sec. III-B), and (3) a high-level vision-language-based planning and execution monitoring module (Sec. III-C). The following sections detail each component.
@@ III-B2 Imitation Learning Module
For the IL policies, we adopt the Humanoid Imitation Transformer (HIT) [10], designed specifically for high-DoF humanoids with binocular vision. It is based on the Action Chunking Transformer (ACT) [31] model, which has demonstrated improved robotic control by predicting multi-step future actions in one forward pass. At inference, HIT outputs action chunks of 50 joint-angle targets at 25 Hz. During training, an auxiliary $L_{2}$ loss between predicted and actual future visual embeddings improves visual grounding and generalization. The predicted joint angles are subsequently transformed into corresponding 3D keypoints via forward kinematics, matching the inputs required by the low-level tracking controller.
@@
containing synchronized state-action pairs recorded at 25 Hz. We collect data for the representative task rearrangement involving picking a bag from a table and placing it onto another. Only successful sequences (30 in total) are retained, segmented into separate skill-specific datasets, i.e., picking and placing, for training two policies. IL policy training employs the demonstration data with hyperparameters specified in Table VI. After training, we evaluated independently each skill policy over 30 autonomous trials, which showed success rates of $90\%$ for picking and $83\%$ for placing.
@@ III-C3 Skill Library and PDDL-like Representation
We represent each skill available to the robot using structured, human-readable descriptions that include the ‘preconditions’ and ‘effects’ fields of an operator definition in Planning Domain Definition Language (PDDL) [1]. Each skill description includes a one-sentence summary, explicit natural-language preconditions, effects, and example verification queries, as illustrated for the Pick skill in Fig. 3. Unlike rigid symbolic representations, this PDDL-like form balances logical rigor and flexibility, enabling the planner to perform common-sense reasoning without exhaustive state enumeration.
@@ III-D Full Hierarchical System
The complete integrated system implements the hierarchical pipeline presented in Fig. 2. The pipeline comprises: (1) a high-level VLM planner (GPT-4o) that generates a sequence of skills from initial visual inputs and natural-language task instructions, (2) a real-time VLM skill monitor (Gemini-2.0-Flash-Lite) that verifies skill execution status based on visual feedback, (3) a skill selector that dynamically chooses the current skill to execute based on the planned sequence and monitor feedback, (4) a mid-level imitation learning Module that activates the corresponding imitation-learned policy according to the selected skill, producing joint-motion targets from egocentric images and proprioceptive data, and (5) a low-level tracking policy combined with a high-frequency PD controller, converting these joint targets into executable motor torque commands.
@@ IV-C Experimental Procedure
A total of 40 independent trials were conducted. The robot operated fully autonomously, executing plans generated by the high-level VLM planner, with skills dynamically selected and monitored for completion in real-time by the VLM execution monitor. Success was defined by the robot correctly picking up the bag and successfully placing it on the designated target surface.
@@ IV-D Experimental Results
TABLE I: Evaluation of individual policies and the full system.
Pick Place Pick-Place
Number of trials 30 30 40
Number of successes 27 25 29
Success rate 90% 83% 73%
Over 40 trials, the integrated hierarchical system successfully completed the entire manipulation sequence with an overall success rate of $73\%$, as summarized in Table I. During each trial, we recorded failures at three distinct levels of the hierarchical framework. The observed failures were categorized as follows, in order of frequency:
@@ V Discussion
The experiments demonstrate that augmenting the canonical two-layer humanoid stack with a vision-language planning and monitoring layer enables reliable execution of multi-step non-prehensile manipulation routines in the physical world. Deployed on a 29-DoF Unitree G1, our proposed three-level architecture successfully completed the obstructed bag-transfer task in 73% of forty independent trials. The execution monitor’s one-second mean verification latency at 1Hz fits comfortably within the robot’s motion cycle, and coexists with high-frequency skill policies and RL controller, illustrating that the large-model reasoning and perception need not preclude high-rate control.
Compared to prior humanoid manipulation systems, which often rely on open-loop sequences of predefined skills, our approach introduces continuous closed-loop reasoning. By integrating interpretable VLM-based planning with real-time visual verification, our framework dynamically selects appropriate skills and explicitly determines when transitions between skills should occur.
@@ V-A Failures and Limitations
A qualitative failure analysis reveals three dominant modes. First, skill-policy drift occurs when partial occlusions or out-of-distribution object poses cause the imitation policy to stall short of a grasp or placement. Second, planner misgrounding emerges when the deliberative VLM hallucinates objects or fails to sequence skills correctly. This typically involves adding additional irrelevant plan steps or formulating redundant transition questions. Third, optimistic monitor judgments occasionally trigger a premature transition when subtle geometric conditions—such as a bag resting on the edge of a table—are misclassified as success.
Several limitations temper the present findings. Our chosen evaluation task family, while non-trivial, is confined to table-top pick-and-place scenes and therefore under-samples the combinatorial diversity of domestic manipulation. The 1Hz verification rate, although adequate for deliberative motions, would be insufficient for fast skills, such as rapid recovery motions. This implicitly defines the timescale of skill abstraction.
@@ V-B Future Work
Error detection and adaptive re-planning:  A critical future enhancement includes adapting the VLM execution monitor system to explicitly detect execution failures or anomalous states [23, 26]. Through failure detection, we can trigger real-time adaptive replanning. This promises to improve the robustness of the system.
Enhanced semantic and contextual conditioning:  Future work involves incorporating richer conditioning beyond visual and proprioceptive states from the high-level execution monitor. For example, replacing the ResNet backbone with that of CLIP [24] would allow the VLM to have finer-grained semantic control of the skills via text-conditioning.
Expanding the skill library:  Finally, the current skill library is focused primarily on pick-and-place. An immediate next step involves expanding the robot’s skill repertoire. Examples include non-prehensile bi-manipulation and tool use. This broader skill set would enable more complex, realistic long-horizon tasks, requiring the planning and monitoring components to handle increasingly diverse and dynamic environments.
@@ VI Conclusion
In this work, we presented a hierarchical planning and control framework for multi-step humanoid manipulation tasks, combining robust RL-based whole-body tracking, imitation-learned skill policies, and a high-level vision-language planning and monitoring module. Evaluated on the Unitree G1 humanoid robot, our integrated system demonstrated successful autonomous execution of complex manipulation sequences, achieving a 73% success rate across real-world trials. Crucially, our approach offers inherent modularity, interpretability, and adaptability—enabling straightforward extension of the robot’s skill repertoire. This work highlights the effectiveness of integrating vision-language models with hierarchical control, paving the way toward increasingly capable autonomous humanoid systems operating in dynamic human environments.